julia is 8 years older then dylan. sofia is four years youner than dylan. the sum of there ages is 55. what is each persons age?
Answer by ankor@dixie-net.com(22740)

julia is 8 years older then dylan.
j = d + 8
sofia is four years younger than dylan.
s = d - 4
the sum of there ages is 55.
j + d + s = 55
what is each persons age?
replace j with (d+8), replace s with (d-4)
(d+8) + d + (d-4) = 55
combine like terms
3d + 4 = 55
3d = 55 - 4
3d = 51
d = 51/3
d = 17 yrs is Dylan's age
:
I'm sure you can find the other ages, ensure they all add up to 55